Richard took a loan of $10,000 at a rate of 7%. If he paid $1,400 in interest, how many years did it take him to pay back the loan?
Answer by mananth(16949)

Rate, r = 7%=0.07
Loan, p =10,000
Interest, i = 1400
i=p*r*t
1400=10000*0.07*t
1400/(0.07*10000)=t
t= 2 years